# Topological restrictions on relatively Anosov representations

### v1

@@ 5.2. Excluding spheres in the boundary
$\mathsf{R}(x)=\underbrace{\mathsf{R}_{1}(x)}_{\in\xi_{\rho^{\ast}}^{d-k}(x_{0})}+\underbrace{\mathsf{R}_{2}(x)}_{\in\xi_{\rho}^{k}(x_{0})}$ (4)
where $\mathsf{R}_{1}\colon S^{m}\to\xi_{\rho^{\ast}}^{d-k}(x_{0})$ and $\mathsf{R}_{2}\colon S^{m}\to\xi_{\rho}^{k}(x_{0})$ are continuous functions. In particular, we obtain the decomposition for the map $\widetilde{\mathsf{H}}$:
$\widetilde{\mathsf{H}}(x)=\mathsf{R}(x)+\underbrace{\mathsf{G}_{0}(-x)}_{\in\xi_{\rho}^{k}(x_{0})}+\underbrace{v_{1}}_{\in\xi_{\rho^{\ast}}^{d-k}(x_{0})}=\underbrace{\mathsf{R}_{1}(x)+v_{1}}_{\in\xi_{\rho^{\ast}}^{d-k}(x_{0})}+\underbrace{\mathsf{R}_{2}(x)+\mathsf{G}_{0}(-x)}_{\in\xi_{\rho}^{k}(x_{0})}\quad\forall\ x\in S^{m}.$ (5)
@@
$\widetilde{\mathsf{F}}(x)=\underbrace{\mathsf{F}_{0}(x)}_{\in\xi_{\rho^{\ast}}^{d-k}(x_{0})}+\underbrace{\omega_{1}}_{\in\xi_{\rho}^{k}(x_{0})}\quad\forall\ x\in S^{m}.$ (6)
Since $\widetilde{\mathsf{H}}(x)\in\xi_{\rho}^{d-k}(f(x))\cap\xi_{\rho}^{d-k}(f(-x))$, by (1), we have that the vector $\widetilde{\mathsf{F}}(\pm x)\in\mathbb{R}^{d}$ is perpendicular to $\widetilde{\mathsf{H}}(x)\in\mathbb{R}^{d}$ for every $x\in S^{m}$. In other words, by (3) and (5), (6) we have that
@@
We remark that, by (5), $\mathsf{R}_{1}(x)+v_{1}\neq 0$ since $\widetilde{\mathsf{H}}(x)\in\xi_{\rho}^{d-k}(f(x))\subset\mathbb{R}^{d}\smallsetminus\xi_{\rho}^{k}(x_{0})$. Also, the map $\mathsf{F}_{0}\colon S^{m}\to\xi_{\rho^{\ast}}^{d-k}(x_{0})$ is injective since $\xi_{\rho^{\ast}}^{k}$ is transverse to itself and $\mathsf{F}$ is an injective map. Finally, we obtain the equality
@@
where $\mathsf{F}_{0}^{\prime}\colon S^{m}\to S^{1}(\xi_{\rho^{\ast}}^{d-k}(x_{0}))$ is the continuous map
@@
which is odd and hence has non-zero degree. Let us also fix $\sigma\colon S^{1}(\xi_{\rho^{\ast}}^{d-k}(x_{0}))\to S^{m}$ a homeomorphism sending antipodal pairs to antipodal pairs, i.e. $\sigma(-u)=-\sigma(u)$ for every $u\in S^{1}(\xi_{\rho^{\ast}}^{d-k}(x_{0}))$. By using (7) we obtain the following continuous map $A\colon S^{m}\times[-\pi,\pi]\to S^{m}$ defined by
@@
$\Omega(x)=\frac{\mathsf{R}_{1}(x)+v_{1}}{\left\|\mathsf{R}_{1}(x)+v_{1}\right\|}\in S^{1}(\xi_{\rho^{\ast}}^{d-k}(x_{0}))\ \ x\in S^{m}.$
@@
Now there are three cases to consider to finish the proof.
Case 1. Suppose that $d\geq 2k+2,m\geq k+1$. Note that $\textup{dim}(W_{0})\geq 2$, so we may choose a unit vector $v_{2}\in W_{0}:=\xi_{\rho}^{d-k}(x_{0})\cap\xi_{\rho^{\ast}}^{d-k}(x_{0})$ linearly independent from $v_{1}\in W_{0}$. By (a) we see that $\sigma\circ\Omega$ is homotopic to the odd map $\sigma\circ\mathsf{F}_{0}^{\prime}$, so $\sigma\circ\Omega$ has non-zero degree and hence is a surjective onto $S^{m}$. In particular, $\Omega:S^{m}\rightarrow S^{1}(\xi_{\rho^{\ast}}^{d-k}(x_{0}))$ is surjective and there is $y_{0}\in S^{m}$ such that
@@
Note that $\mathsf{R}(y_{0})\neq{\bf 0}$ since $\lambda_{0}v_{2}-v_{1}\in\xi_{\rho^{\ast}}^{d-k}(x_{0})\smallsetminus\{\bf 0\}$ and $\mathsf{R}_{2}(w_{0})\in\xi_{\rho}^{k}(x_{0})$. On the other hand, since $\lambda_{0}v_{2}-v_{1}\in\xi_{\rho^{\ast}}^{d-k}(x_{0})\cap\xi_{\rho}^{d-k}(x_{0})$, we have $\mathsf{R}(y_{0})\in\xi_{\rho}^{d-k}(x_{0})$. However, this contradicts the fact that $\mathsf{R}(x)\in\xi_{\rho}^{k}(f(-x))$ for every $x\in S^{m}$, and $\xi_{\rho}^{k}(f(-y_{0}))\cap\xi_{\rho}^{d-k}(x_{0})=\{\bf 0\}$. The theorem follows in this case.
Case 2. Suppose that $d=2k+1,m=k$ and $k\in\mathbb{N}$ is even. Recall that $\sigma\circ F_{0}^{\prime}\colon S^{k}\to S^{k}$ is odd and has non-zero degree. By (b), since $\sigma\circ\mathsf{F}_{0}^{\prime}$ and $\theta\circ\sigma\circ\mathsf{F}_{0}^{\prime}$ are homotopic and the antipodal map $\theta:S^{k}\rightarrow S^{k}$ has degree $-1$, we have that $\textup{deg}(\sigma\circ\mathsf{F}_{0}^{\prime})=\textup{deg}(\theta\circ\sigma\circ F_{0}^{\prime})=-\textup{deg}(\sigma\circ\mathsf{F}_{0}^{\prime})$, so $\textup{deg}(\sigma\circ\mathsf{F}_{0}^{\prime})=0$. This is a contradiction, and hence the theorem follows in this case.
Case 3. Suppose that $d=2k+1,m=k$ and $k\in\mathbb{N}$ is odd. Note that since $x_{0}\in\partial(\Gamma,P)\smallsetminus f(S^{k})$, by transversality, the map $\tau_{k}^{+}\circ\xi_{\rho}^{k}\circ f\colon S^{k}\to\mathbb{P}(\wedge^{k}\operatorname{\mathbb{R}}^{d})$ avoids the hyperplane $\tau_{k}^{-}(\xi^{k+1}_{\rho}(x_{0}))\subset\wedge^{k}\mathbb{R}^{2k+1}$. Thus, we obtain a continuous lift $\widetilde{\xi}_{k}\colon S^{k}\to S^{1}(\wedge^{k}\operatorname{\mathbb{R}}^{d})$ of $\tau_{k}^{+}\circ\xi_{\rho}^{k}\circ f$. In addition, the continuous map $\mathsf{F}_{1}\colon S^{k}\to\mathbb{P}(\xi_{\rho}^{k+1}(x_{0}))\smallsetminus\mathbb{P}(\xi_{\rho}^{k}(x_{0}))$ defined by
$\mathsf{F}_{1}(x)=\big{[}\xi_{\rho}^{k+1}(f(x))\cap\xi_{\rho}^{k+1}(x_{0})\big{]}$
admits a continuous lift, i.e. there is $v_{0}\in\xi_{\rho}^{k+1}(x_{0})\smallsetminus\xi_{\rho}^{k}(x_{0})$ and a continuous map $\mathsf{F}_{1}^{\prime}\colon S^{k}\to\xi_{\rho}^{k}(x_{0})\cong\operatorname{\mathbb{R}}^{k}$ such that $[\mathsf{F}_{1}^{\prime}(x)+v_{0}]=[\mathsf{F}_{1}(x)]$ for every $x\in S^{k}$.
By transversality for $x,y\in S^{k}$ distinct we have that $\xi_{\rho}^{k+1}(f(x))=\xi_{\rho}^{k}(f(x))\oplus\mathbb{R}\mathsf{F}_{1}(x)$ and
$\operatorname{\mathbb{R}}^{d}=\xi_{\rho}^{k}(f(x))\oplus\xi_{\rho}^{k}(f(y))\oplus\mathbb{R}\mathsf{F}_{1}(x),$
so we obtain the continuous, non-vanishing map $\mathcal{H}\colon(S^{k})^{(2)}\to\operatorname{\mathbb{R}}^{\ast}$ defined as follows
Now the map $\mathsf{F}_{1}^{\prime}$ cannot be injective so there are $x_{1},y_{1}\in S^{k}$ distinct with $\mathsf{F}_{1}^{\prime}(x_{1})=\mathsf{F}_{1}^{\prime}(y_{1})$. But then, since $k\in\mathbb{N}$ is odd, we have
$=-\mathcal{H}(x_{1},y_{1}).$
This contradicts the fact that $\operatorname{\mathcal{H}}$ is a non-zero continuous map. This finishes the proof of Case 3. The proof of the theorem is complete.∎
By adapting the argument in Case 3 of the proof of Theorem 5.3, we obtain a positive answer to Question 7.1 in the case of $k$-Anosov representations into $\mathsf{SL}_{2k+1}(\mathbb{R})$ when $k\in\mathbb{N}$ is odd and the Gromov boundary $\partial_{\infty}\Delta$ is connected.
@@
The group $\Delta$ is one-ended, hence its Gromov boundary $\partial_{\infty}\Delta$ is locally connected and has no global cut points by [Bow98]. Since $\partial_{\infty}\Delta$ is not the circle, we may find distinct points $z,w\in\partial_{\infty}\Delta$, such that $\partial_{\infty}\Delta\smallsetminus\{z,w\}$ is connected. In particular, the set $X(z)$ is non-empty.
@@
$\delta^{+},\delta^{-}\in U\smallsetminus\{z,w,z^{-},w^{-}\},\ g\delta^{+}\neq\delta^{-}\ \forall g\in F.$
@@
Observe that since $\delta^{q}z,\delta^{2q}w\in U\smallsetminus\{z^{-},w^{-}\}$, we have $\lim_{n\rightarrow\infty}x_{n}=z$ and $\lim_{n\rightarrow-\infty}x_{n}=w$. Since $\partial_{\infty}\Delta\smallsetminus\{z,w\}$ is connected, [Bar18, Lem. 2.5] implies that there exists $p\in\mathbb{Z}$ such that $\partial_{\infty}\Delta\smallsetminus\{x_{p},x_{p+1}\}$ is connected.
@@
By transversality, the image of the map $\tau_{k}^{+}\circ\xi_{\rho}^{k}:\partial_{\infty}\Delta\smallsetminus\{z\}\rightarrow\mathbb{P}(\wedge^{k}\mathbb{R}^{2k+1})$ avoids the hyperplane $\mathbb{P}(\tau_{k}^{-}(\xi_{\rho}^{k+1}(z)))$ and hence lifts to a continuous map $\widetilde{\xi}_{k}:\partial_{\infty}\Delta\smallsetminus\{z\}\rightarrow S^{1}(\wedge^{k}\mathbb{R}^{2k+1})$. Thus, the map $\mathcal{H}_{z}\colon(\partial_{\infty}\Delta\smallsetminus\{z\})^{(2)}\rightarrow\mathbb{R}$ given by
@@
Claim. $\mathcal{H}_{z}(x,y)\ \mathcal{H}_{z}(y,x)>0$ for every $x\neq y$.
Let $x,y\in\partial_{\infty}\Delta\smallsetminus\{z\}$ and $U_{x},U_{y}\subset\partial_{\infty}\Delta\smallsetminus\{z\}$ be arbitrary disjoint open neighbourhoods containing $x,y\neq z$ respectively. Since $\partial_{\infty}\Delta$ is locally connected [Bow98], we may assume that $U_{x},U_{y}$ are connected. Since $X(z)$ is dense in $\partial_{\infty}\Delta$ we may choose $x^{\prime}\in U_{x}\cap X(z)$, $x^{\prime\prime}\in U_{x}\cap X(z)$, $x^{\prime}\neq x^{\prime\prime}$, and $y^{\prime}\in U_{y}\cap X(z)$. Now observe that $(x,y)$ and $(x^{\prime},y^{\prime})$ lie in the connected subset $U_{x}\times U_{y}$ of $(\partial_{\infty}\Delta\smallsetminus\{z\})^{(2)}$. The pairs $(x^{\prime},y^{\prime})$ and $(x^{\prime},x^{\prime\prime})$, lie in the connected subset $\{x^{\prime}\}\times(\partial_{\infty}\Delta\smallsetminus\{z,x^{\prime}\})$, and $(x^{\prime},x^{\prime\prime})$ and $(y^{\prime},x^{\prime\prime})$ lie in the connected set $(\partial_{\infty}\Delta\smallsetminus\{z,x^{\prime\prime}\})\times\{x^{\prime\prime}\}$. We conclude that, since $\mathcal{H}_{z}$ is continuous, we have $\mathcal{H}_{z}(x,y)\mathcal{H}_{z}(y,x^{\prime\prime})>0$. Since $U_{x},U_{y}$ are arbitrary open neighbourhoods and $\mathcal{H}_{z}(x,y),\mathcal{H}_{z}(y,x)\neq 0$, we have $\mathcal{H}_{z}(x,y)\mathcal{H}_{z}(y,x)>0$ for every $x\neq y$ and the claim follows.
@@ 6. Constraints in high cohomological dimension
Let $d\geq 4$. If $\rho\colon(\Gamma,\operatorname{\mathcal{P}})\to\operatorname{\mathsf{SL}}_{d}(\operatorname{\mathbb{R}})$ is a relatively 1-Anosov representation, and $\Gamma$ has cohomological dimension at least $d-1$, then $\operatorname{\mathcal{P}}=\varnothing$, $\Gamma$ is hyperbolic and $\rho$ is a Benoist representation.
The proof of Theorem 6.1 follows the proof of [CT20, Thm. 1.5]
@@
First, we assume that $\partial(\Gamma,\mathcal{P})$ is connected. Note that since $\Gamma$ cannot be a surface group, by Theorem 2.2, $\partial(\Gamma,\mathcal{P})$ is not homeomorphic to $S^{1}$. Let $V:=\langle\xi_{\rho}^{1}(\partial(\Gamma,\mathcal{P}))\rangle$. By Propositions 2.3 and 2.12(1) the limit set $\xi_{\rho}^{1}(\partial(\Gamma,\mathcal{P}))$ is contained in an affine chart $A$ of $\mathbb{P}(V)$. Note that since $\Gamma$ acts as a convergence group on $\partial(\Gamma,\mathcal{P})$, the restriction $\rho|_{V}$ has to have finite kernel and discrete image by Observation 3.1. In particular, the image $\rho|_{V}(\Gamma)$ preserves the convex hull $\mathcal{C}_{\rho}$ of $\xi_{\rho}^{1}(\partial(\Gamma,\mathcal{P}))$ in $A$ and its interior $\Omega_{\rho}:=\textup{Int}(\mathcal{C}_{\rho})$. By assumption our assumption that $\textup{cd}(\Gamma)\geq d-1$, it follows that $\rho|_{V}(\Gamma)$ acts cocompactly on $\Omega_{\rho}$, $\rho(\Gamma)$ cannot contain unipotents and $V=\mathbb{R}^{d}$. In other words, $\mathcal{P}=\varnothing$, $\Gamma$ is word hyperbolic, $\rho$ is $1$-Anosov and $\partial\Omega_{\rho}=\xi_{\rho}^{1}(\partial_{\infty}\Gamma)$. Moreover, by Benoist’s theorem [Ben04] the domain $\Omega_{\rho}$ is strictly convex.
It remains to rule out the case where $\partial(\Gamma,\mathcal{P})$ is disconnected. Indeed, if this is the case, since $\Gamma$ is torsion-free, by Theorem 2.7, there is a splitting $\Gamma=\Gamma_{1}\ast\cdots\ast\Gamma_{r}$ and finite families of subgroups $\mathcal{P}_{i}$ of $\Gamma_{i}$ such that $\partial(\Gamma_{i},\mathcal{P}_{i})$ is connected for every $1\leq i\leq r$. Since $\textup{cd}(\Gamma)=\max_{i}\textup{cd}(\Gamma_{i})$, we may assume that $\textup{cd}(\Gamma_{1})\geq d-1$. By the previous case, we deduce that $\mathcal{P}_{1}=\varnothing$ and there is a properly convex domain $\Omega_{1}\subset\mathbb{P}(\mathbb{R}^{d})$ preserved and acted cocompactly by $\rho(\Gamma_{1})$. Then $\rho|_{\Gamma_{1}}$ is $1$-Anosov and

### v2

@@ 5.2. Excluding spheres in the boundary
$\mathsf{R}(x)=\underbrace{\mathsf{R}_{1}(x)}_{\in(\pi_{k}(x_{0}))^{\perp}}+\underbrace{\mathsf{R}_{2}(x)}_{\in\pi_{k}(x_{0})}$ (3)
where $\mathsf{R}_{1}\colon S^{m}\to(\pi_{k}(x_{0}))^{\perp}$ and $\mathsf{R}_{2}\colon S^{m}\to\pi_{k}(x_{0})$ are continuous functions. In particular, we obtain the decomposition for the map $\widetilde{\mathsf{H}}$:
$\widetilde{\mathsf{H}}(x)=\mathsf{R}(x)+\underbrace{\mathsf{G}_{0}(-x)}_{\in\pi_{k}(x_{0})}+\underbrace{v_{1}}_{\in(\pi_{k}(x_{0}))^{\perp}}=\underbrace{\mathsf{R}_{1}(x)+v_{1}}_{\in(\pi_{k}(x_{0}))^{\perp}}+\underbrace{\mathsf{R}_{2}(x)+\mathsf{G}_{0}(-x)}_{\in\pi_{k}(x_{0})}\quad\forall\ x\in S^{m}.$ (4)
@@
$\widetilde{\mathsf{F}}(x)=\underbrace{\mathsf{F}_{0}(x)}_{\in(\pi_{k}(x_{0}))^{\perp}}+\underbrace{\omega_{1}}_{\in\pi_{k}(x_{0})}\quad\forall\ x\in S^{m}.$ (5)
Since $\widetilde{\mathsf{H}}(x)\in\pi_{d-k}(f(x))\cap\pi_{d-k}(f(-x))$, we have that the vector $\widetilde{\mathsf{F}}(\pm x)\in\mathbb{R}^{d}$ is perpendicular to $\widetilde{\mathsf{H}}(x)\in\mathbb{R}^{d}$ for every $x\in S^{m}$. In other words, by (2) and (4), (5) we have that
@@
We remark that, by (4), $\mathsf{R}_{1}(x)+v_{1}\neq 0$ since $\widetilde{\mathsf{H}}(x)\in\pi_{d-k}(f(x))\subset\mathbb{R}^{d}\smallsetminus\pi_{k}(x_{0})$. Also, the map $\mathsf{F}_{0}\colon S^{m}\to(\pi_{k}(x_{0}))^{\perp}$ is injective since $\pi_{k}$ and hence $\mathsf{F}$ is injective. Finally, we obtain the equality
@@
where $\mathsf{F}_{0}^{\prime}\colon S^{m}\to\mathbb{S}((\pi_{k}(x_{0}))^{\perp})$ is the continuous map
@@
which is odd and hence has non-zero degree. Let us also fix $\sigma\colon\mathbb{S}((\pi_{k}(x_{0}))^{\perp})\to S^{m}$ a homeomorphism sending antipodal pairs to antipodal pairs, i.e. $\sigma(-u)=-\sigma(u)$ for every $u\in\mathbb{S}((\pi_{k}(x_{0}))^{\perp})$. By using (6) we obtain the following continuous map $A\colon S^{m}\times[-\pi,\pi]\to S^{m}$ defined by
@@
$\Omega(x)=\frac{\mathsf{R}_{1}(x)+v_{1}}{\left\|\mathsf{R}_{1}(x)+v_{1}\right\|}\in\mathbb{S}(\xi_{\rho^{\ast}}^{d-k}(x_{0}))\ \ x\in S^{m}.$
@@
Now there are two cases to consider to finish the proof.
Case 1. Suppose that $d\geq 2k+2$ (so $m\geq k+1$). Note that $\textup{dim}(W_{0})\geq 2$, so we may choose a unit vector $v_{2}\in W_{0}:=\pi_{d-k}(x_{0})\cap(\pi_{k}(x_{0}))^{\perp}$ linearly independent from $v_{1}\in W_{0}$. By (a) we see that $\sigma\circ\Omega$ is homotopic to the odd map $\sigma\circ\mathsf{F}_{0}^{\prime}$, so $\sigma\circ\Omega$ has non-zero degree and hence is a surjective onto $S^{m}$. In particular, $\Omega:S^{m}\rightarrow\mathbb{S}((\pi_{k}(x_{0}))^{\perp})$ is surjective and there is $y_{0}\in S^{m}$ such that
@@
Note that $\mathsf{R}(y_{0})\neq{\bf 0}$ since $\lambda_{0}v_{2}-v_{1}\in(\pi_{k}(x_{0}))^{\perp}\smallsetminus\{\bf 0\}$ and $\mathsf{R}_{2}(w_{0})\in\pi_{k}(x_{0})$. On the other hand, since $\lambda_{0}v_{2}-v_{1}\in(\pi_{k}(x_{0}))^{\perp}\cap\pi_{d-k}(x_{0})$, we have $\mathsf{R}(y_{0})\in\pi_{d-k}(x_{0})$. However, this contradicts the fact that $\mathsf{R}(x)\in\pi_{k}(f(-x))$ for every $x\in S^{m}$, and $\pi_{k}(f(-y_{0}))\cap\pi_{d-k}(x_{0})=\{\bf 0\}$.
Case 2. Suppose that $d=2k+1$ (so $m=k$) and $k\in\mathbb{N}$ is even. Recall that $\sigma\circ F_{0}^{\prime}\colon S^{k}\to S^{k}$ is odd and has non-zero degree. By (b), since $\sigma\circ\mathsf{F}_{0}^{\prime}$ and $\theta\circ\sigma\circ\mathsf{F}_{0}^{\prime}$ are homotopic and the antipodal map $\theta:S^{k}\rightarrow S^{k}$ has degree $-1$, we have that $\textup{deg}(\sigma\circ\mathsf{F}_{0}^{\prime})=\textup{deg}(\theta\circ\sigma\circ F_{0}^{\prime})=-\textup{deg}(\sigma\circ\mathsf{F}_{0}^{\prime})$, so $\textup{deg}(\sigma\circ\mathsf{F}_{0}^{\prime})=0$. This is a contradiction.
In either case we obtain a contradiction, and hence the proof of the theorem is complete.∎
By adapting the argument in Lemma 5.4, we obtain a positive answer to Question 7.1 in the case of $k$-Anosov representations into $\mathsf{SL}_{2k+1}(\mathbb{R})$ when $k\in\mathbb{N}$ is odd and the Gromov boundary $\partial_{\infty}\Delta$ is connected.
@@
The group $\Delta$ is one-ended, hence its Gromov boundary $\partial_{\infty}\Delta$ is locally connected and has no global cut points by [Bowditch-connected]. Since $\partial_{\infty}\Delta$ is not the circle, we may find distinct points $z,w\in\partial_{\infty}\Delta$, such that $\partial_{\infty}\Delta\smallsetminus\{z,w\}$ is connected. In particular, the set $X(z)$ is non-empty.
@@
$\delta^{+},\delta^{-}\in U\smallsetminus\{z,w,z^{-},w^{-}\},\ g\delta^{+}\neq\delta^{\pm}\ \forall g\in F.$
@@
Observe that since $\delta^{q}z,\delta^{2q}w\in U\smallsetminus\{z^{-},w^{-}\}$, we have $\lim_{n\rightarrow\infty}x_{n}=z$ and $\lim_{n\rightarrow-\infty}x_{n}=w$. Since $\partial_{\infty}\Delta\smallsetminus\{z,w\}$ is connected, [Barrett, Lem. 2.5] implies that there exists $p\in\mathbb{Z}$ such that $\partial_{\infty}\Delta\smallsetminus\{x_{p},x_{p+1}\}$ is connected.
@@
By transversality, the image of the map $\tau_{k}^{+}\circ\xi_{\rho}^{k}:\partial_{\infty}\Delta\smallsetminus\{z\}\rightarrow\mathbb{P}(\wedge^{k}\mathbb{R}^{2k+1})$ avoids the hyperplane $\mathbb{P}(\tau_{k}^{-}(\xi_{\rho}^{k+1}(z)))$ and hence lifts to a continuous map $\widetilde{\xi}_{k}:\partial_{\infty}\Delta\smallsetminus\{z\}\rightarrow\mathbb{S}(\wedge^{k}\mathbb{R}^{2k+1})$. Thus, the map $\mathcal{H}_{z}\colon(\partial_{\infty}\Delta\smallsetminus\{z\})^{(2)}\rightarrow\mathbb{R}$ given by
@@
Claim: $\mathcal{H}_{z}(x,y)\ \mathcal{H}_{z}(y,x)>0$ for every $x\neq y$.
Let $x,y\in\partial_{\infty}\Delta\smallsetminus\{z\}$ and $U_{x},U_{y}\subset\partial_{\infty}\Delta\smallsetminus\{z\}$ be arbitrary disjoint open neighbourhoods containing $x,y\neq z$ respectively. Since $\partial_{\infty}\Delta$ is locally connected [Bowditch-connected], we may assume that $U_{x},U_{y}$ are connected. Since $X(z)$ is dense in $\partial_{\infty}\Delta$ we may choose $x^{\prime}\in U_{x}\cap X(z)$, $x^{\prime\prime}\in U_{x}\cap X(z)$, $x^{\prime}\neq x^{\prime\prime}$, and $y^{\prime}\in U_{y}\cap X(z)$. Now observe that $(x,y)$ and $(x^{\prime},y^{\prime})$ lie in the connected subset $U_{x}\times U_{y}$ of $(\partial_{\infty}\Delta\smallsetminus\{z\})^{(2)}$. The pairs $(x^{\prime},y^{\prime})$ and $(x^{\prime},x^{\prime\prime})$, lie in the connected subset $\{x^{\prime}\}\times(\partial_{\infty}\Delta\smallsetminus\{z,x^{\prime}\})$, and $(x^{\prime},x^{\prime\prime})$ and $(y^{\prime},x^{\prime\prime})$ lie in the connected set $(\partial_{\infty}\Delta\smallsetminus\{z,x^{\prime\prime}\})\times\{x^{\prime\prime}\}$. Finally, $(y^{\prime},x^{\prime\prime})$ and $(y,x^{\prime\prime})$ lie in the connected set $U_{y}\times\{x^{\prime\prime}\}$, and $(y,x^{\prime\prime})$ and $(y,x)$ lie in the connected set $\{y\}\times U_{x}$. We conclude that, since $\mathcal{H}_{z}$ is continuous, and $\mathcal{H}_{z}(x,y),\mathcal{H}_{z}(y,x)\neq 0$, we have $\mathcal{H}_{z}(x,y)\mathcal{H}_{z}(y,x)>0$ for every $x\neq y$. ∎
@@ 6. Constraints in high cohomological dimension
Let $d\geq 4$. If $\rho\colon(\Gamma,\operatorname{\mathcal{P}})\to\operatorname{\mathsf{SL}}_{d}(\operatorname{\mathbb{R}})$ is a relatively 1-Anosov representation, and $\Gamma$ is a torsion-free group of cohomological dimension at least $d-1$, then $\operatorname{\mathcal{P}}=\varnothing$, $\Gamma$ is hyperbolic and $\rho$ is a Benoist representation.
The proof of Theorem 6.1 follows the proof of [CanaryTsouvalas, Thm. 1.5]
@@
First, we assume that $\partial(\Gamma,\mathcal{P})$ is connected. Note that since $\Gamma$ cannot be a free group or surface group, $\partial(\Gamma,\mathcal{P})$ is not homeomorphic to $S^{1}$ by Theorem 2.1. Let $V:=\langle\xi_{\rho}^{1}(\partial(\Gamma,\mathcal{P}))\rangle$. By Propositions 2.3 and 2.12(1) the limit set $\xi_{\rho}^{1}(\partial(\Gamma,\mathcal{P}))$ is contained in an affine chart $A$ of $\mathbb{P}(V)$. By Proposition 2.12(3), $\rho(\Gamma)$ preserves a properly convex domain $\Omega_{\rho}\subset\mathbb{P}(\operatorname{\mathbb{R}}^{d})$. By our assumption that $\textup{cd}(\Gamma)\geq d-1$, it follows that $\rho|_{V}(\Gamma)$ acts cocompactly on $\Omega_{\rho}$. Hence $\rho(\Gamma)$ cannot contain elements with all eigenvalues of modulus one [DGK, Prop. 10.3] and $V=\mathbb{R}^{d}$. Since, for a relatively Anosov representation, all the eigenvalues of the image of a peripheral element have modulus one [ZZ1, Prop. 4.2], we conclude here we must have $\mathcal{P}=\varnothing$. Hence $\Gamma$ is word hyperbolic, $\rho$ is $1$-Anosov and (by cocompactness) $\partial\Omega_{\rho}=\xi_{\rho}^{1}(\partial_{\infty}\Gamma)$. Moreover, by Benoist’s theorem [BenoistI] the domain $\Omega_{\rho}$ is strictly convex.
It remains to rule out the case where $\partial(\Gamma,\mathcal{P})$ is disconnected. Indeed, if this is the case, since $\Gamma$ is torsion-free, by Theorem 2.7, there is a splitting $\Gamma=\Gamma_{1}\ast\cdots\ast\Gamma_{r}$ and finite families of subgroups $\mathcal{P}_{i}$ of $\Gamma_{i}$ such that $\partial(\Gamma_{i},\mathcal{P}_{i})$ is connected for every $1\leq i\leq r$. Since $\textup{cd}(\Gamma)=\max_{i}\textup{cd}(\Gamma_{i})$, we may assume that $\textup{cd}(\Gamma_{1})\geq d-1$. By the previous case, we deduce that $\mathcal{P}_{1}=\varnothing$ and there is a properly convex domain $\Omega_{1}\subset\mathbb{P}(\mathbb{R}^{d})$ such that $\rho(\Gamma_{1})$ preserves and acts cocompactly on $\Omega_{1}$. Then $\rho|_{\Gamma_{1}}$ is $1$-Anosov and